OpenStudy (anonymous):

what is the slope of the line 4x+6y=12

OpenStudy (anonymous):

wouldn't it be -2/3

OpenStudy (anonymous):

You have to put it in this form: \(y=mx+b\) So: 6y=-4x+12 y=-2/3x+2 slope is m

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw